what is the square root of 8x to the power of 3?
Answer by nyc_function(2741)

sqrt[8x^3] =
House into individual radicals and then simplify.
sqrt[8] = sqrt[4] * sqrt[2]
sqrt[4] * sqrt[2] = 2(sqrt[2])
==============================
sqrt[x^2] * sqrt[x] = x(sqrt[x])
==============================
We now have the following:
2(sqrt[2]) * x(sqrt[x]) =
2x(sqrt[2x])